I need help with this story problem.
A triangular lot has a perimeter of 54 yards of fencing. The longest side of the lot is 11 yards longer than the shortest side. The other side is 7 yards longer than the shortest side. Find the value of the three sides.

Answer by monika_p(71)

Perimeter is a sum of the lenght of all the sides of the triangle
lets say that x is a lenght of the shortest side,
then the longest side y that is 11 yards longer than x we can write as y=x+11. The Other side z is 7 yards longer than x then,
then z=x+7
Perimeter formula P=x+y+z now substitute for P,y, and z so u have one unknown x
54=x+(x+11)+(x+7)
when you solve for x you will get x=12, then y=23, and z =19
Check P=12+23+19=54 true!!!